OpenStudy (anonymous):

log√3 (9) and log1/4 (16) can someone help

OpenStudy (owlcoffee):

let's look at the first one: \[\log_{\sqrt{3}} 9\] If you want to know what's the result of it, we just have to apply the logarithm definition that states: \[\log_{a} b=c <=>a ^{c}=b\] Let's apply it: \[\log_{\sqrt{3}} 9 <=> (\sqrt{3})^{x}=9\] And I'll rewrite the sqrt of 3 and convert the 9 into 3^2: \[(3^{\frac{ 1 }{ 2 }})^{x}=3^{2}\] And since there the exponent of the exponent, I just multiply them: \[3^{\frac{ x }{ 2 }}=3^{2}\] The base of that exponential equation is the same on both sides, so I can work with the exponents only: \[\frac{ x }{ 2 }=2\] Solving for x we conclude that x=4. So we have finished, we say that: \[\log_{\sqrt{3}} 9=4\]

OpenStudy (yttrium):

SWAG!!!

OpenStudy (owlcoffee):

swag!

OpenStudy (anonymous):

Awesome, thank you Owlcoffee.

OpenStudy (owlcoffee):

Try applying that to the other one.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I think it will be -2

OpenStudy (owlcoffee):

that's correct.

OpenStudy (anonymous):

thank's